The first winner in a lottery get #2000 while subsequent winners get #50 less than the preceeding winner. how much does the 20th position get?
Answer by mananth(16946)


The first winner in a lottery get #2000 while subsequent winners get #50 less than the preceeding winner. how much does the 20th position get?
a=2000
d=-50
t20 =?
tn = a+(n-1)d
t20 = 2000+(19)(-50)
t20 = 2000 -950
t20 = 1050
the 20th position gets 1050